Operations Report – January 20th

Crew 203 Operations Report 20-JAN-2019

SOL: 8

Name of person filing report: Freddy Castaneda

Non-nominal systems: Nothing new to report.

Notes on non-nominal systems: None

Generator (hours run): 13hr 13min; Turned on last night (19Jan2019) at 19:33; Turned off this morning (20Jan2019) at 8:46; Turned on tonight (20Jan2019) at 17:40

Solar SOC – Turned on (19Jan2019) 60%; Turned off (20Jan2019) 60%; Turned on (20Jan2019) 86%

Diesel Reading – Just under 1/2 tank.

Propane Reading – 21%

Ethanol Free Gasoline – 10 gallons were added.

Water (auxiliary tank) – Not in use

Water (static tank) – Over 45%; About 240 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – Over 59%; 178 gallons

Water (loft) – At level marker 12

Static to Loft Pump used – Yes; At 9:45 and 18:00 to refill tank

Water Meter: 01403017

Toilet tank emptied: Yes.

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Not used

Hours: 68.8

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Opportunity rover used: Used to test brakes.

Hours: 46.0

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Not used

Hours: 70.6

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Brakes on Opportunity were tested but still not working properly.

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: None

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? Used to take trash to Janksville.

CrewCar used and why, where? Off campus

General notes and comments: Nothing to report.

Summary of internet: Nothing to report.

Summary of suits and radios: Suit #3 still pending for review.

Summary of Hab operations: Nothing to report.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: Nothing to report